the pressure of a quantity of an ideal gas was held constant. the initial volume and volume and temperature were 20 liters and 273 K, respectively. what would be the final temperature if the final volume was 30 liters?
Answer by scott8148(6628)

at constant pressure , the volume and absolute temperature are directly related

20 / 30 = 273 / t

t = (30 * 273) / 20